# Towards Graph Foundation Models: A Survey and Beyond

### v3

@@ 1 Introduction
Much like the evolution witnessed in NLP, graph machine learning is also undergoing a paradigm transition. In its early stages, graph tasks predominantly employed shallow methods, such as random walk [6] and matrix factorization [7]. These methods, however, were typically limited to transductive learning on unattributed graphs [8]. The more recent shift towards deep learning methods has catalyzed the rise of graph neural networks (GNNs). GNNs have revolutionized the landscape by introducing the message-passing mechanism, where nodes iteratively aggregate information from their neighbors. By harnessing GNNs in fully supervised, semi-supervised, or unsupervised settings, researchers have pioneered a variety of customized graph models. These advancements have yielded substantial improvements in tasks like node classification [9], link prediction [10], graph classification [11], and graph clustering [12]. However, certain challenges of GNN models still persist. For example, GNNs are restricted with issues related to expressive power [13] and generalizability [14], especially given the ever-expanding datasets and the widening spectrum of tasks.
@@
Inspired by the success of LLMs as foundation models in NLP, researchers have explored the possibilities of graph foundation models towards the emergence and homogenization capabilities. These explorations primarily focus on the design of backbone architectures for GFMs, and different learning paradigms including pre-training and adaptation, as they are the key strategies of LLMs to acheive the aforementioned capabilities. First and foremost, the emergent abilities of foundation models typically exist only in backbones with a large number of parameters, whereas the parameter count of graph neural networks is significantly smaller than that of the language backbones. This implies that the backbone of graph foundation models may need to be redesigned to achieve more substantial knowledge storage towards emergence. As graph data is typically associated with rich text information, an alternative approach is to use LLMs as graph foundation models. Nonetheless, it remains uncertain whether LLMs can effectively handle graph data and associated tasks, and it is crucial to determine how to model graph structures in LLMs. Additionally, the homogenization of foundation models necessitates the handling of diverse tasks in a uniform manner. Devising effective pretext tasks and downstream task adaptation methods are challenging for graph data, due to the complexity in interconnected nodes and various forms of attributes, as well as the diversity in tasks across node-, edge- and graph-levels. Therefore, there is also a need to design suitable pre-training tasks and adaptation mechanisms.
While there is no definitive solution for designing and implementing graph foundation models, this paper surveys some related researches and categorizes them into three distinct approaches based on their reliance on GNNs and LLMs: (1) GNN-based Models: They aim to enhance existing graph learning paradigms through innovations in the backbone, pre-training, and adaptation aspects; (2) LLM-based Models: They explore the feasibility of using an LLM as a graph foundation model by converting graphs into text or tokens; (3) GNN+LLM-based Models: They explore various forms of synergy between GNNs and LLMs to empower them with enhanced capabilities.
@@
  • This article defines the concept of graph foundation models for the first time, and examines the core issues and characteristics of their capabilities.
  • This article introduces a novel taxonomy and discusses the strengths and limitations of each approach towards graph foundation models.
  • This article provides promising future directions towards graph foundation models.
The subsequent sections of this article are organized as follows. In Section 2, we introduce the background related to graph foundation models. Section 3 defines graph foundation models and highlights their similarities and differences with language foundation models. Sections 4 - 6 delve into the relevant works that consider GNN-based models, LLM-based models and GNN+LLM-based models as graph foundation models, separately. Section 7 engages in a discussion on the future directions of graph foundation models. In Section 8, we summarize the key points of this paper.
@@ 2 Background
Before introducing the concept of graph foundation models, in this section, we first review some background knowledge, namely, deep graph learning and language foundation models. Specifically, we introduce them from three aspects: data, backbone architectures, and learning paradigms.
@@ 2.1 Deep Graph Learning
Many real-world systems naturally find their representations in the form of graphs. Deep graph learning holds significant importance across various fields for its powerfulness in modeling and representing complex relationships and interactions among entities. In this section, we provide a concise overview that covers the main steps of deep graph learning, which consists of three key aspects: graph data, backbone architectures and learning paradigms.
@@ 2.1.1 Graph Data
Graph is a versatile and powerful data representation that captures intricate relationships and dependencies among entities in a network. The graph data has the following several characteristics. (1) Non-Euclidean Nature: Graph data is inherently non-Euclidean, as it lacks the rigid geometric structure found in traditional data formats [19]. In contrast to Euclidean data with fixed neighborhood region and a deterministic order, graph data explicitly encode complex connections between entities. (2) Various Domains: Graph data is ubiquitous across various domains, including social networks [20], biology [21] and transportation [22], etc. Graph data in various domains can exhibit distinct characteristics, including different node types, edge semantics, and structural patterns. For example, in biological networks, nodes can represent proteins, genes, or metabolites, while edges may indicate interactions such as protein-protein interactions or metabolic reactions. This domain-specific variability makes it challenging to create a one-size-fits-all model that can effectively generalize and adapt to diverse graph structures [23]. (3) Various Types: Graph data comes in various types, including homogeneous graphs [24], heterogeneous graphs [25], hypergraphs [26] and dynamic graphs [27], etc. Homogeneous graphs contain nodes and edges of the same type, such as a citation network composed of papers. Heterogeneous graphs include more than one type of node or edge, such as a citation network containing both authors and papers. Hypergraphs are composed of hyperedges connecting multiple nodes and can model high-order relationships among nodes. Dynamic graphs refer to graph structures where nodes and edges change over time, as seen in a traffic network formed by changing traffic flow.
@@ 2.1.2 Backbone Architectures
As the current mainstream backbone architecture, graph neural networks (GNNs) have emerged as a powerful framework for deep graph learning. Most GNNs follow the message-passing framework [28], which enables nodes in a graph to exchange information with their neighbors. For example, GCN [9] introduces the concept of graph convolution layers and lays the foundation for many subsequent GNN architectures. GraphSAGE [29] proposes a method for generating embeddings for nodes in large graphs using inductive learning. Additionally, GAT [30] introduces the attention mechanism to GNNs, allowing nodes to weigh the importance of their neighbors during message passing, enhancing their expressive power. These works have significantly contributed to the advancement of GNNs, making them versatile tools for deep graph learning.
Despite the fact that deeper neural networks can achieve stronger expressive power [31], deepening GNNs is not easy. The reason is that as the number of layers in GNN increases, too much information is introduced into the message aggregation process, causing the representations of all nodes to become similar [32]. This is also known as the over-smoothing problem. Furthermore, an increase in the number of layers leads to an exponential growth in the receptive field, giving rise to the over-squashing problem [13], where a substantial amount of information is compressed into fixed-length node vectors. In recent years, several efforts have been made to address the over-smoothing problem and over-squashing problem in deep graph neural networks, resulting in improved downstream task performance. For example, innovations like DropEdge [33] enhance GCN models by randomly removing edges for improved performance and scalability (up to 64 layers). Another thread of work to improve the expressive power of GNNs is graph transformer [34, 35, 36]. Thanks to its fully connected attention mechanism and the long-range relationship modeling capability, the graph transformer architecture can alleviate the over-smoothing problem and over-squashing problem [37].
@@ 2.1.3 Learning Paradigms
The learning paradigms for deep graph learning encompass three primary categories: supervised learning, semi-supervised learning and unsupervised learning. In this section, we will provide a brief introduction to these learning paradigms.
Supervised learning. In the supervised setting, algorithms leverage a training dataset comprising input data paired with corresponding output labels. This paradigm finds practical applications in tasks such as graph classification [38] and graph regression [39]. For example, in the molecular property prediction task [40], GNN models are trained to predict specific chemical properties or attributes of molecules using labeled training data, enabling the discovery of valuable insights for drug development and materials research.
Semi-supervised learning. Semi-supervised learning, as highlighted in a recent study [41], constitutes the primary focus of deep graph learning. This approach harnesses both labeled and unlabeled data to enhance model performance, with node classification [9] emerging as a prominent application. The message-passing mechanism [28] empowers GNNs to iteratively exchange information among neighboring nodes. This capability enables the propagation of information throughout the graph, effectively incorporating both labeled and unlabeled data to facilitate predictions. Furthermore, GNNs can also be combined with traditional methods like label propagation to further enhance their performance in semi-supervised settings [42].
Unsupervised learning. Unsupervised learning [43] is a broader machine learning approach to learn patterns and structures from data without manual labels. For example, graph clustering [44] aims to discover inherent structures and patterns within the graph solely based on the relationships and connectivity between nodes. Another example is link prediction, which aims to predict missing or upcoming connection relationships. An important subclass of unsupervised learning is self-supervised learning, which aims to generate labels using information inherent in the data itself [45]. Based on self-supervised learning, GNNs can be trained end-to-end and applied to downstream tasks such as graph clustering [12] and link prediction [10].
@@ 2.2 Language Foundation Models
AI is currently undergoing a transformative shift marked by the emergence of some specific natural language models (such as GPT-3) that are trained on extensive and diverse datasets using large-scale self-supervised learning. These models, known as foundation models, are able to produce a wide array of outputs, enabling them to tackle a broad spectrum of downstream tasks. In contrast to the deep graph learning pipeline, the foundation model’s approach embraces a pre-training and adaptation framework, enabling the model to achieve several significant advancements, including the emergence [2] and homogenization [1]. Foundation models have primarily established themselves in the field of NLP [1], so our discussion will focus on language foundation models in this section.
@@ 2.2.1 Language Data
Language data refers to text or spoken content in a human language, encompassing the grammar rules of the natural language and the associated semantics of the words. It can include written documents, transcribed audio recordings, and any other form of language-based communication. Language data is essential for many NLP tasks, such as machine translation, sentiment analysis and text summarization. Researchers and developers use language data to train and evaluate language models and other NLP algorithms. The quality and quantity of language data play a crucial role in the performance of NLP systems, impacting their accuracy, robustness, and overall effectiveness in various language tasks. In contrast to computer vision and other domains, the size of annotated language data is rather small, consisting of only a few thousand sentences [46]. This limitation is primarily due to the high cost of manual annotation. Nevertheless, there is a vast amount of unlabeled language data available from sources such as the internet, newspapers, and books, creating opportunities for utilizing unlabeled data in model pre-training. Furthermore, compared to graph data, language data as Euclidean data is easier to model, and its rich semantic information significantly enhances the knowledge transferability of language models.
@@ 2.2.2 Backbone Architectures
An early breakthrough in foundation models is pre-trained language models (PLMs), designed to capture context-aware word representations, which proved remarkably effective as versatile semantic features. For instance, BERT [47], grounded in the parallelizable Transformer architecture [48] with self-attention mechanisms, is conceived through the pre-training of bidirectional language models with specifically designed pretext tasks on vast unlabeled data. This landmark study significantly elevates the performance benchmarks for NLP tasks and serves as a catalyst for a plethora of subsequent research, establishing the prevailing pre-training and fine-tuning learning paradigm.
Furthermore, researchers have observed that increasing the scale of PLMs, whether by augmenting model size or training data, frequently results in increased model capacity for downstream tasks. These larger PLMs, collectively referred to as LLMs, exhibit distinctive behaviors compared to their smaller counterparts (e.g., the 1.5B-parameter GPT-2 and 175B-parameter GPT-3). After training on massive text datasets, they manifest remarkable capabilities, often referred to as emergent abilities [2], such as in-context learning [1]. LLMs primarily utilize the Transformer architecture, because highly parallelizable Transformer-based architectures accelerate the pre-training stage and enable the utilization of massive datasets. In the context of Transformer models, tokens serve as the input and represent units at the word level in natural language texts. Typically, LLMs containing hundreds of billions (or more) of parameters [49], exemplified by models such as GPT-3 [50], PaLM [51], Galactica [52], and LLaMA [53].
@@ 2.2.3 Learning Paradigms
Pre-train and Fine-tune. In this paradigm, a model with a consistent architecture is initially pre-trained as a language model (LM), where it predicts the probability of observed textual data. In comparison to end-to-end training, pre-training offers distinct advantages and forms cornerstone for the capabilities of foundation models. Firstly, pre-training on huge text corpus enables the learning of universal language representations, which is a possible reason to explain emergent abilities [54]. Additionally, pre-training provides improved model initialization, typically resulting in enhanced generalization performance, enabling homogenization of multiple tasks [54]. Furthermore, pre-training serves as a form of regularization, helping to prevent overfitting on smaller datasets [56]. For instance, models like GPT-3 [50] are trained with a language modeling objective, rewarding them for their abilities in emergence and homogenization.
@@
Following the pre-training phase, foundation models acquire general-purpose capabilities suitable for a broad spectrum of tasks. Nevertheless, pre-trained models still lack downstream task-specific information, and using them directly may not yield optimal results. Therefore, we need to tune the model for specific tasks, which is known as fine-tuning. Building upon the success of models like ULMFit [57] and BERT [47], fine-tuning has emerged as the predominant method for adapting pre-trained models. In this framework, the primary emphasis lies in objective engineering, encompassing the design of training objectives for both pre-training and fine-tuning phases. For instance, Pegasus [58] shows that incorporating a loss function for predicting important sentences within a document results in an improved pre-trained model for text summarization. The advantage of fine-tuning is that it can transfer knowledge between source and target tasks (or domains) and benefit the model’s performance. For the small size of fine-tuning dataset compared to pre-training dataset, this process can enable adaptation effectively without losing the stored structural language knowledge.
Pre-train, Prompt and Predict. In this paradigm, rather than adjusting pre-trained language models to suit specific downstream tasks, the approach involves reshaping the downstream tasks to align more closely with those tackled during the original LM training, accomplished by providing textual prompts. By selecting suitable prompts, we can steer the LM’s behavior so that it can predict the desired output, sometimes without any additional task-specific training. This method offers the advantage of enabling a single, entirely unsupervised LM, when equipped with a set of fitting prompts, to handle a wide array of tasks [55].
@@
From the aspect of prompt engineering, the approaches to create a proper prompts can be classified to manual methods and automated methods. Manual methods involve creating intuitive templates based on human insight, which is the most straightforward approach to crafting prompts. For instance, the influential LAMA dataset [59] offers manually devised cloze templates to assess the knowledge of language models. However, manual methods face challenges in terms of high cost and precision. To address these issues, some approaches have started to experiment with automated prompt generation. For example, Prompt Mining [60] is a template discovery approach that autonomously identifies templates based on a given set of training inputs and outputs.
Looking at it from a different perspective, in terms of how models and prompts are combined to generate results, prompting strategies can be categorized into three approaches: tuning-free prompting, prompt tuning, and instruction tuning [5]. Tuning-free prompting directly generates answers based solely on a prompt without altering the parameters of the pre-trained LLMs [53] [61]. Prompt tuning introduces supplementary prompt-relevant parameters in addition to the parameters of the pre-trained models and update these additional parameters using supervision signals acquired from downstream training samples [62] [63]. Instruction tuning adapts the LM’s parameters in a manner similar to the fine-tuning process, while additionally incorporating fixed instructions to guide the model’s behavior. This approach offers potential enhancements, especially in zero-shot scenarios [64].
@@ 3 Graph foundation models
In this section, we will first formally define the concepts of graph foundation models, including the definition, key characteristics and key technologies. Then, we will discuss the impact from graph data and graph tasks on graph foundation models. Finally, we will discuss the similarities and differences between graph foundation models and language foundation models.
3.1 Concepts of Graph Foundation Model
In this subsection, we will first provide a definition of graph foundation models. Following that, we will delve into the key characteristics and essential techniques of graph foundation models, as well as the impact of graph data and graph tasks on graph foundation models.
3.1.1 Definition and Key Characteristics
We define a graph foundation model as follows:
Emergence. Emergence means that the graph foundation model will exhibit some new abilities when having a large parameters or trained on more data, which are also referred to as emergent abilities. Inspired by the various emergent abilities [67, 68, 69] possessed by foundation models, we expect GFMs to have similar abilities, including in-context learning, graph reasoning, and zero-shot graph generation, etc. In-context learning allows predictions for various downstream tasks with few examples [70]. Graph reasoning decomposes a complex problem into multiple sub-problems based on the graph structure and addresses them step by step, such as solving graph algorithm problems [71]. Zero-shot graph generation requires the model to generate graphs based on the desired conditions without the need for any examples [72]. Note that although language foundation models have demonstrated various emergent abilities, only a few works [70, 71, 72] have explored emergent abilities of GFMs so far.
3.1.2 Key Technologies
3.1.3 Impact from Graph Data
Graph Type. Based on the number of categories of nodes and edges in a graph, we can categorize graphs into homogeneous graphs and heterogeneous graphs. In homogeneous graphs, all nodes and edges belong to the same category. For example, in a social graph where nodes represent individuals (users) and edges represent friendship relationships, it is a homogeneous graph because all nodes are individuals and all edges represent friendship relationships. Heterogeneous graphs, on the other hand, have more than one type of nodes or edges, representing different types of entities and relationships [25]. For instance, an e-commerce graph may include nodes for users, products, and purchase relationships, forming a heterogeneous graph. For graph foundation models, handling heterogeneous graphs poses greater challenges and typically requires the design of specific backbone architectures and optimization objectives. Nonetheless, utilizing the meta-path based approach [80], a heterogeneous graph can be mapped into multiple homogeneous graphs, one for each meta-path. For example, one can apply the GFM trained on homogeneous graphs to each of these meta-path induced homogeneous graphs, separately, to get node embedding. Then, these embeddings on homogeneous graphs under different meta-paths can be fused together. However, beyond homogeneous graphs and heterogeneous graphs, there are some more complex types of graphs in the real world, such as dynamic graphs and hypergraphs [81], which poses additional challenges for GFM.
@@ 3.1.3 Impact from Graph Data
3.1.4 Impact from Graph Tasks
Node-level Tasks. Node-level tasks refer to the classification, regression, or prediction performed on each node. Common node-level tasks include node classification, node regression, and clustering coefficient prediction. For example, in social networks, graph nodes can represent users, and node classification can be used to identify users from different social circles or with varying interests.
In summary, the format of tasks in graphs are highly diverse and can be categorized into three types: node-level, edge-level, and graph-level, each of which has wide-ranging applications. This undoubtedly increases the challenge of homogenization for graph foundation models. For example, in graph classification and node classification tasks on synthetic datasets, modeling structural information is often more crucial [87]. On the other hand, when dealing with node classification tasks on graphs with rich node features, modeling feature information becomes more important [87]. Furthermore, tasks that are more similar to each other will also have a lower transfer difficulty, implying that these tasks are more likely to be addressed using the same GFM. While increasing expressive power holds promise for improving the performance of many node-level, edge-level, and graph-level tasks [88], there is also some work suggesting that overly strong expressive power may not be necessary for graph generation tasks [89].
3.2 Comparison with Language Foundation Model

### v4

# Graph Foundation Models: Concepts, Opportunities and Challenges

@@ 1 Introduction
Much like the evolution witnessed in NLP, graph machine learning is also undergoing a paradigm transition. In its early stages, graph tasks predominantly employed shallow methods, such as random walk [6, 7] and matrix factorization [8, 9, 10, 11, 12]. These methods, however, were typically limited to transductive learning [13]. The more recent shift towards deep learning methods has catalyzed the rise of graph neural networks (GNNs). GNNs have revolutionized the landscape by introducing the message-passing mechanism, where nodes iteratively aggregate information from their neighbors. By harnessing GNNs in fully supervised, semi-supervised, or unsupervised settings, researchers have pioneered a variety of customized graph models. These advancements have yielded substantial improvements in tasks like node classification [14], link prediction [15], graph classification [16], and graph clustering [17]. However, certain challenges of GNN models still persist. For example, GNNs are restricted with issues related to expressive power [18] and generalizability [19], especially given the ever-expanding datasets and the widening spectrum of tasks.
@@
Inspired by the success of LLMs as foundation models in NLP, researchers have explored the possibilities of graph foundation models towards the emergence and homogenization capabilities. These explorations primarily focus on the design of backbone architectures for GFMs, and different learning paradigms including pre-training and adaptation, as they are the key strategies of LLMs to acheive the aforementioned capabilities. First and foremost, the emergent abilities of foundation models typically exist only in backbones with a large number of parameters, whereas the parameter count of GNNs is significantly smaller than that of the language backbones. This implies that the backbone of GFMs may need to be redesigned to achieve more substantial knowledge storage towards emergence. As graph data is typically associated with rich text information, an alternative approach is to use LLMs as GFMs. Nonetheless, it remains uncertain whether LLMs can effectively handle graph data and associated tasks, and it is crucial to determine how to model graph structures in LLMs. Additionally, the homogenization of foundation models necessitates the handling of diverse tasks in a uniform manner. Devising effective pre-training tasks (also called pretext tasks) and downstream task adaptation methods are challenging for graph data, due to the complexity in interconnected nodes and various forms of attributes, as well as the diversity in tasks across node-, edge- and graph-levels. Therefore, there is also a need to design suitable pre-training tasks and adaptation mechanisms.
While there is no definitive solution for designing and implementing GFMs, this paper surveys some related researches and categorizes them into three distinct approaches based on their reliance on GNNs and LLMs: (1) GNN-based Models: They aim to enhance existing graph learning paradigms through innovations in the backbone, pre-training, and adaptation aspects; (2) LLM-based Models: They explore the feasibility of using an LLM as a GFM by converting graphs into text or tokens; (3) GNN+LLM-based Models: They explore various forms of synergy between GNNs and LLMs to empower them with enhanced capabilities.
@@
$\bullet$ This article defines the concept of graph foundation models for the first time, and examines the core issues and characteristics of their capabilities.
$\bullet$ This article introduces a novel taxonomy and discusses the strengths and limitations of each approach towards graph foundation models.
$\bullet$ This article provides promising future directions towards graph foundation models.
The subsequent sections are organized as follows. In Section 2, we introduce the background related to GFMs. Section 3 defines GFMs and highlights their similarities and differences with language foundation models. Sections 4 - 6 delve into the relevant works that consider GNN-based models, LLM-based models and GNN+LLM-based models as GFMs, separately. Section 7 engages in a discussion on the future directions of GFMs. In Section 8, we summarize the key points of this paper.
@@ 2 Background
Before introducing GFMs, we review background knowledge on deep graph learning and language foundation models.
@@ 2.1 Deep Graph Learning
This section provides a concise overview from three key aspects: graph data, backbone architectures, and learning paradigms.
@@ 2.1.1 Graph Data
Graphs capture intricate relationships among entities and possess several key characteristics that make them challenging for machine learning tasks. The primary challenge stems from their (1) Non-Euclidean Nature: Graph data is inherently non-Euclidean [24], lacking the rigid geometric structure of traditional data formats such as 1D text, 2D images or tabular data. This means that graph data cannot be adequately described within a simple flat space because its intrinsic structure does not conform to the principles of Euclidean geometry [25]. Unlike Euclidean data, which often comes in predefined shapes (e.g., images of a specific resolution), non-Euclidean data can vary greatly in size and shape. This variability complicates the design of algorithms that often navigate complex topologies, significantly increasing computational cost compared to operations on simpler Euclidean spaces.
Beyond this fundamental structural complexity, two additional challenges are posed by the nature of graph data. (2) Various Domains: Graph data appears in domains such as social networks [26], biology [27], and transportation [28]. It is also used in tasks like 3D human skeleton recognition [29], semantic segmentation [30], and video classification [31]. Domain-specific variability with different node types and edge semantics makes creating a universal model challenging [32]. (3) Various Types: graph data includes homogeneous, heterogeneous [33], hyper- [34], and dynamic ones [35]. Such diversity also brings challenges to deep graph learning.
@@ 2.1.2 Backbone Architectures
GNNs are the current mainstream backbone architecture for deep graph learning. Most GNNs follow the message-passing framework [36], enabling nodes to exchange information with neighbors. For example, GCN [14] introduces graph convolution layers, GraphSAGE [37] generates node embeddings using inductive learning, and GAT [38] adds an attention mechanism to weigh neighbor importance, enhancing expressive power. These contributions make GNNs versatile tools for deep graph learning.
However, deepening GNNs is challenging. Increasing layers leads to over-smoothing, where node representations become similar [39], and over-squashing, where information is overly compressed [18]. Efforts like DropEdge [40], which randomly removes edges, improve GCN performance and scalability. Graph transformers [41, 42, 43], with their fully connected attention and long-range relationship modeling, help alleviate over-smoothing and over-squashing [44].
@@ 2.1.3 Learning Paradigms
The learning paradigms for deep graph learning encompass three primary categories:
Supervised learning. In supervised learning, algorithms use a training dataset with input data and output labels. This is used in tasks like graph classification [45] and graph regression [46]. For instance, in molecular property prediction [47], GNNs predict chemical properties using labeled data, aiding drug development and materials research.
Semi-supervised learning. Semi-supervised learning, as discussed in [48], is a primary focus in deep graph learning. It uses both labeled and unlabeled data to improve model performance, with node classification [14] being a key application. The message-passing mechanism [36] allows GNNs to exchange information among nodes, incorporating both data types for predictions. GNNs can also combine with methods like label propagation for better performance [49].
Unsupervised learning. Unsupervised learning discovers patterns and structures without manual labels. Graph clustering identifies structures based on node relationships, while link prediction estimates missing connections. A subclass, self-supervised learning, generates labels from the data itself [50], allowing GNNs to be trained end-to-end for tasks like graph clustering [17] and link prediction [15].
@@ 2.2 Language Foundation Models
AI is currently undergoing a transformative shift marked by the emergence of some specific language models (such as GPT-3) that are trained on extensive and diverse datasets using self-supervised learning. These models, known as foundation models, are able to produce a wide array of outputs, enabling them to tackle a broad spectrum of downstream tasks. In contrast to the deep graph learning pipeline, the foundation model’s approach embraces a pre-training and adaptation framework, enabling the model to achieve several significant advancements, including the emergence [2] and homogenization [1]. Foundation models have primarily established themselves in the field of NLP [1], so our discussion will focus on language foundation models in this section.
@@ 2.2.1 Language Data
Language data refers to text or spoken content in a human language, encompassing the grammar rules of the natural language and the associated semantics of the words. The quality and quantity of language data play a crucial role in the performance of NLP systems, impacting their accuracy, robustness, and overall effectiveness in various language tasks. In contrast to graph data, language data as Euclidean data is easier to model, and its rich semantic information significantly enhances the knowledge transferability of language models.
@@ 2.2.2 Backbone Architectures
An early breakthrough in foundation models is pre-trained language models (PLMs), designed to capture context-aware word representations, which proved remarkably effective as versatile semantic features. Furthermore, researchers have observed that increasing the scale of PLMs, whether by augmenting model size or training data, frequently results in increased model capacity for downstream tasks. These larger PLMs, collectively referred to as LLMs, exhibit emergent abilities [2] compared to their smaller counterparts (e.g., the 1.5B-parameter GPT-2 and 175B-parameter GPT-3). LLMs primarily utilize the Transformer architecture, because highly parallelizable Transformer-based architectures accelerate the pre-training stage and enable the utilization of massive datasets. In Transformer models, tokens serve as the input and represent units at the word level in natural language texts.
@@ 2.2.3 Learning Paradigms
Pre-train and Fine-tune. In this paradigm, a model is initially pre-trained as a language model (LM), where it predicts the probability of observed textual data. Following the pre-training phase, we need to tune the model for specific tasks, which is known as fine-tuning. Building upon the success of models like ULMFit [53] and BERT [54], fine-tuning has emerged as the predominant method for adapting pre-trained models. In this framework, the primary emphasis lies in objective engineering, encompassing the design of training objectives for both pre-training and fine-tuning phases. The advantage of fine-tuning is that it can transfer knowledge between source and target tasks (or domains) and benefit the model’s performance. For the small size of fine-tuning dataset compared to pre-training dataset, this process can enable adaptation effectively without losing much pre-trained knowledge.
@@
Pre-train, Prompt and Predict. In this paradigm, rather than adjusting PLMs to suit specific downstream tasks, the approach involves reshaping the downstream tasks to align more closely with those tackled during the original LM training, accomplished by providing textual prompts. From the aspect of prompt engineering, the approaches to create a proper prompts can be classified to manual methods and automated methods. Manual methods involve creating intuitive templates based on human insight, which is the most straightforward approach to crafting prompts. Manual methods face challenges in terms of high cost and precision. To address these issues, some approaches have started to experiment with automated prompt generation. When LLMs have billions of parameters, it is more effective if we can just adapt the downstream tasks without adjusting model parameters. This makes prompt-tuning a promising approach for adapting LLMs.
@@ 3 Graph foundation models
In this section, we will first formally define the concepts of graph foundation models, including the definition, key characteristics and key technologies. Then, we will discuss the similarities and differences between graph and language foundation models.
3.1 Definition and Key Characteristics
We define a graph foundation model as follows:
Emergence. Emergence means that the graph foundation model will exhibit some new abilities when having a large parameters or trained on more data, which are also referred to as emergent abilities. Inspired by the various emergent abilities [57, 58] possessed by foundation models, we expect GFMs to have similar abilities, including in-context learning, graph reasoning, and zero-shot graph generation, etc. In-context learning allows predictions for various downstream tasks with few examples [59]. Graph reasoning decomposes a complex problem into multiple sub-problems based on the graph structure and addresses them step by step, such as solving graph algorithm problems [60]. Zero-shot graph generation requires the model to generate graphs based on the desired conditions without the need for any examples [61]. Note that although language foundation models have demonstrated various emergent abilities, only a few works [59, 60, 61] have explored emergent abilities of GFMs so far.
3.2 Key Technologies
3.3 Comparison between GFMs and LLMs